OpenStudy (anonymous):

1. think of a number but tell it . 2. add the same number in it. 3. add 40 in it 4. divide the total by 2 5. subtract the remaining by the original number 6. the answer you'll get is 20 !!!

OpenStudy (anonymous):

@SkoobyLoveYou @rebeccaskell94 @lgbasallote check it out

OpenStudy (anonymous):

dont tell the number

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

thanks i used to perform this trick when i was in my school

OpenStudy (phi):

1. think of a number but tell it . x 2. add the same number in it. x+ x=2x 3. add 40 in it 2x+40 4. divide the total by 2 (2x+40)/2 = x+20 5. subtract the remaining by the original number x+20-x= 20 6. the answer you'll get is 20 !!! yes

OpenStudy (anonymous):

Let n be the number. 1. n +n =2n 2. 2n + 40 3. (2n+40)/2 = n+20 4 n+20-n = 20 Not interesting :(

OpenStudy (anonymous):

@RolyPoly if u dont think its interesting then dont look at it ...

OpenStudy (anonymous):

Sorry, I meant no offense. But I thought it was a math question.

OpenStudy (anonymous):

Good Job

OpenStudy (anonymous):

thanks skooby

OpenStudy (anonymous):

@RolyPoly no problem but it is a game of maths

OpenStudy (anonymous):

I used 2 do this when I was little I luved doing it was great fun! ;)

OpenStudy (anonymous):

yeah .... its funny stuff

OpenStudy (anonymous):

the number is 0 right?

OpenStudy (anonymous):

which number is zero???

OpenStudy (anonymous):

the number you picked?

OpenStudy (ajprincess):

it's really intersting.

OpenStudy (anonymous):

thanks aj.... really appreciate

OpenStudy (anonymous):

basically i learned it from my papa....

OpenStudy (anonymous):

. nice :)

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

pick any number multiply with 0 you get 0

OpenStudy (ajprincess):

There is another 1 like this. Think of a two digit number. Find the sum of the digits. Subtract that from the original number. The answer u get will be a multiple of 9.

OpenStudy (anonymous):

ok lets mold the trick 1. think of a number but dont tell it . 2. add the same number in it. 3. add 30 in it 4. divide the total by 2 5. subtract the remaining by the original number 6. the answer you'll get is 15 !!!

OpenStudy (anonymous):

here's another

OpenStudy (anonymous):

pick a no 15 the answer u get is 15!!!

OpenStudy (anonymous):

i m telling the exact answer...

OpenStudy (anonymous):

that is the BEAUTY of the trick

OpenStudy (anonymous):

if u pick 15 the answer u'll get is 7.5

OpenStudy (anonymous):

correct!!

OpenStudy (anonymous):

thanks guyz for appreciation

OpenStudy (ajprincess):

yw